# A pipeline for enabling path-specific causal fairness in observational health data

### v1

@@ 4.2. Robustly estimate path-specific causal effects
Despite progress in doubly robust estimators, the dimensionality of $W$ and $Z$ remains a statistical bottleneck for finite sample estimators. We therefore assess the robustness of dimensionality reduction methods that provide the best tradeoff for robust path-specific estimation using synthetic data, where true effect sizes are known. We set the causal graph underlying the data-generating process to the SFM (Figure 1a) and initialize a different MLP to simulate each edge of the graph. We simulate synthetic data with dimensionality that closely matches the real data, with general function mappings (such as highly non-linear MLPs) to ensure that chosen dimensionality reduction will provide analogous robust estimates in real data where the ground-truth is unknown. Synthetic experiments show that doubly robust estimation alone still leads to large measurement errors in high dimensions ($\geq 500$ features) for the NDE and NIE (Figure 5). We empirically evaluate methods of dimensionality reduction to identify the method(s) that lead to the most accurate path-specific effect estimates. Doubly robust estimators are effective because they only require one set of models (either the propensity models or outcome models) to be correctly specified (Li and Shen, 2020). Thus, we focus on dimensionality reduction methods that seek to retain predictive power about the sensitive attribute $X$ or $Y$. For each method listed below, we test reduction to $20\%$, $40\%$, $50\%$, $60\%$, and $80\%$ of the original dimensionality $|W|$.
@@
  • Learn $Y$: We retain predictive power in $Y$ by identifying the ”most important” subset of mediating features $W$ for predicting the outcome $Y$. We examine importance through LASSO logistic regression, commonly used for genetic mediation analysis (Jérolon et al., 2024; Ye et al., 2021). Additionally, we train an XGBoost-based outcome model and identify features through permutation feature importance (PFI) (Fisher et al., 2019), a model-agnostic measure of feature importance.
@@
4.3. Train a baseline model
The baseline model is trained on all variables in $X$, $W$, and $Z$. Our pipeline is model-agnostic, so any model architecture can be used. The evaluation of causal effects on model output requires ”hard” (binary) labels; we create these labels by selecting the threshold probability that jointly optimizes sensitivity and specificity in the validation dataset.
4.4. Select target effects for which to reduce disparities
Once we select a clinical problem and establish a baseline model, we should select the causal paths to prioritize for elimination. This idea of prioritizing causal pathways is similar to the notion of ”business necessity” (Plečko and Bareinboim, 2024); however, rather than ignoring bias through specific pathways, we recognize that regularization across all causal pathways may not be feasible while maintaining model performance (”no free lunch”). Thus, we advocate for case-specific prioritization of causal pathways that accounts for domain knowledge of health disparities and relative disparities in the data compared to baseline models. We examine 1) the relative sizes of each causal effect in the raw data and the ways in each observed causal effect may be reflective of known healthcare disparities in the specific clinical domain; 2) how causal effects may be exacerbated in the baseline model compared to the raw data; and 3) the contextual factors (either from health disparities literature or algorithmic fairness literature) that may explain the causal effects of the baseline model.
4.5. Perform and evaluate fairness interventions
@@
  • Demographic unawareness: We train models without any explicit demographic information (race and gender), including the sensitive attribute. This experiment is representative of the idea of ”fairness through unawareness”, a simple but widely debated debiasing method in fair ML and ML for health (Höltgen and Oliver, 2025; Basu, 2023).
  • ”Unbiased” feature selection: We identify ”highly biased features” as those with a large standardized mean difference (SMD) between the specified sensitive attribute and remove these features during model training. For a specified feature count ($n$), we keep the $n$ features with the lowest SMD.
  • Greedy feature selection We greedily select features that maximize predictive performance while minimizing bias. We quantify a feature’s contribution to predictive performance using a feature importance metric and quantify its contribution to bias using SMD, as defined above. The relative contribution of SMD to the feature score is selected based on the ”elbow point” that maximizes distance from the baseline importance-SMD tradeoff.
@@
  • Equalized Odds: We investigate the impact of imposing a group fairness constraint on path-specific causal fairness. Specifically, we operationalize equalized odds (Hardt et al., 2016) by adding a regularization term to the loss function that penalizes differences in the true and false positive rates.
@@
We evaluate the models based on performance (AUROC) and path-specific causal effects (focusing on the ”prioritized” causal pathways, as described in Section 4.4). In order to better characterize the impact of the regularized prediction model on specific subgroups of patients, we also report the Pearson Correlation between the sensitive attribute and the true outcome, as well as between the sensitive attribute and the predicted outcomes (Appendix D-G). This allows us to compare the relationship between the sensitive attribute and the outcome in the data, baseline model, and the ”fair” model. This analysis is especially important for tasks where our goal is to improve racial fairness, as we report the causal effects with respect to Black vs. White patients, and assessing the impact on other racial groups is crucial.
5. Experiments and Results
@@
We demonstrate our proposed workflow on four clinical risk prediction tasks: acute myocardial infarction (AMI), systemic lupus erythematosus (SLE), type 2 diabetes mellitus (T2DM), and schizophrenia (SCZ). We select these tasks because they span a diverse set of clinical domains and are all associated with known health disparities (Section 5.1). In each case, we select an ”at-risk” cohort of individuals with related symptoms or comorbidities of the illness, and seek to predict which of these individuals will go on to develop the given disease. Our framing of tasks within the ”at-risk” cohort rather than a general population increases the task difficulty because many well-established risk factors for the disease are widely present in the data and are therefore less predictive of the outcome of interest. This forces the model to rely less on risk factors that clinicians are already attuned to, thus increasing potential for clinical utility. For AMI, SLE, and T2DM, we use the phenotypes defined in a prior foundation model benchmark Pang et al. (2025) and retrain the LLAMA foundation model on CUMC-EHR. We select the LLAMA model because is a widely used architecture used both within and outside the healthcare setting (Wornow et al., 2025).
5.1. Define the task and key variables
@@
We demonstrate our pipeline in two datasets representing typical structured observational health data sources: CUMC-EHR, a deidentified EHR of $\sim$5.3 million patients from a large urban academic medical center, and the deidentified Merative Multi-State Medicaid dataset (MDCD). The MDCD dataset includes all data collected from healthcare providers for medical billing for $25$ million patients from $11$ different US states. Both datasets are structured using the Observational Medical Outcomes Partnership Common Data Model (Hripcsak et al., 2019), ensuring reproducibility in other compatible health datasets. In both datasets, gender is recorded as male or female and the method of capture for this variable (e.g., patient-reported, assigned at birth) is not recorded and may vary between patients. We are therefore unable to extend our analyses beyond binary gender. In CUMC-EHR, race is captured as a categorical variable: American Indian or Alaskan Native, Asian, Black, Native Hawaiian or Pacific Islander, White, Missing, and Other. In MDCD, race is captured with three categories: Black, White, and Missing. The method of capture for race is also unknown in both datasets.
5.1.1. Acute Myocardial Infarction (AMI)
@@
We predict AMI in patients with prior ischemic heart disease. We focus on mitigating gender disparities ($x_{0}=female,x_{1}=male$) due to known gender-based differences in presenting symptoms that can lead to underdiagnosis and delayed diagnosis of AMI in women (Kim et al., 2023b; Ngaruiya et al., 2024; Imboden et al., 2026). We predict AMI onset using the embeddings from a LLAMA foundation model (Wornow et al., 2025) pre-trained on CUMC-EHR; the model uses RoPE embeddings to encode temporal information and is trained using next token prediction (NTP) with context length $8192$. The model embeddings represent a person’s clinical history ($W$) and can be optionally combined with $X$ and $Z$ to predict the downstream outcome. This dataset does not contain comprehensive information about visit settings; thus, we represent healthcare utilization as a single variable defined as the number of distinct days with clinical ”touchpoints” (visits, diagnoses, or procedures) per year. $Z$ consists of healthcare utilization and patient race.
5.1.2. Systemic Lupus Erythematosus (SLE)
@@
We predict the first occurrence of SLE among individuals with at least one SLE-related symptom or a prescription of a drug used to treat SLE symptoms. We focus on mitigating gender disparities ($x_{0}=female,x_{1}=male$) in the case of SLE due to differing diagnosis rates ($4-22\%$ of diagnosed SLE patients are male) and significantly differing symptom profiles across sexes (Tan et al., 2012; do Socorro Teixeira Moreira Almeida et al., 2011). Additionally, prior experimental work has found that clinicians faced with otherwise identical symptom profiles are less likely to diagnose men with SLE (Simard et al., 2022). We use the same data source (CUMC-EHR), LLAMA model, and healthcare utilization as outlined above for AMI.
5.1.3. Type 2 Diabetes Mellitus (T2DM)
@@ 5.1.4. Schizophrenia (SCZ)
For the case of SCZ risk prediction, our ”at-risk” cohort is defined as individuals with at least 3 years of observation and at least one psychosis diagnosis, and we predict schizophrenia onset one year after the first psychosis visit. We mandated three years of observation because external validation of the SCZ phenotype demonstrates that requiring three years of prior observation improves the positive predictive value from $69\%$ to $77\%$ (Finnerty et al., 2024). Black individuals are overdiagnosed with schizophrenia due to differences in the interpretation of affective symptoms (Trierweiler et al., 2000; Gara et al., 2019). Additionally, rates of SCZ diagnosis following psychosis are higher for Black individuals compared to white individuals (Moe et al., 2024). For these reasons, we focus on mitigating racial disparities ($x_{0}=White,x_{1}=Black$). The SCZ prediction model is trained for this single task using the MDCD data. We generated features ($W$) based on frequency of conditions, laboratory tests, procedures and duration of prescriptions. We define healthcare utilization through $10$ setting-specific frequency variables (e.g., outpatient, inpatient). $Z$ consists of healthcare utilization and patient gender.
@@
Figure 3. Error in the path-specific causal estimates for dimensionality $|W|=1,750$. Each line corresponds to a different dimensionality reduction method, with the black line corresponding to no dimensionality reduction. The NDE estimation error is reflected in part (a), and the NIE estimation error is reflected in part (b). We find that interventions that identify features important for predicting the outcome ($Y$) outperform all other methods.
@@
Side-by-side line plots with associated error bars. Each line is a different color, corresponding to the given legend of dimensionality reudction techniques
5.2. Robustly estimate path-specific causal effects
@@
Using synthetic data, we demonstrate that estimating causal effects in high mediator dimensions $(|W|\geq 500)$ leads to large estimation errors in the NDE and NIE (Figure 5, Table 2). We test each of the dimensionality reduction methods outlined in Section 4.2 on various mediator dimension sizes, ranging from $|W|=750$ to $|W|=1,750$. We consistently find that dimensionality reduction methods that identify the most important features for predicting $Y$, particularly PFI, are most effective at reducing both the NDE and NIE estimation error across all sample sizes (Figure 3). The spurious effect calculation is not impacted by changing the representation of $W$. These results are broadly consistent for $|W|=750$, $|W|=1,000$ and $|W|=1,500$ (Appendix B). In the case of AMI, SLE, and T2DM, we learn the features most predictive of $Y$ through PFI (Fisher et al., 2019), as this yields the best results in the experiments where $|W|=750$ (and the clinical embeddings are all $|W|=768$). In the case of SCZ, we select highly predictive features using augmented feature occlusion (AFO), a permutation analysis method appropriate for temporal models (Tonekaboni et al., 2020).
5.3. Train a baseline model
@@
Table 1. We report the Natural Direct Effect (NDE), Natural Indirect Effect (NIE), and Spurious Effect (SE) in the dataset and the baseline model for each prediction task. We report the effect size based on the held-out test set to remain consistent with the evaluation of model regularization methods and report accompanying $95\%$ confidence intervals derived through bootstrapping.
5.4. Select target effects for which to reduce disparities
@@
5.4.1. Acute Myocardial Infarction (AMI)
@@
When examining causal effects in the data, only the NIE is significant ($0.008$ $95\%$ CI: $0.005,0.009$]); the positive value of the NIE indicates that an individual with clinical features drawn from a ”male” population distribution is more likely to be diagnosed with AMI. This is consistent with findings that clinicians underdiagnose women with AMI due to a differing ”typical” symptom profile (Kim et al., 2023b; Ngaruiya et al., 2024; Imboden et al., 2026). In the baseline model, both the NDE and NIE increase by an order of magnitude (NDE: $0.073$; NIE: $0.153$), while the SE remains relatively small (0.005[0.008,0.004]-0.005\ [-0.008,\ -0.004]). The increase in NDE and NIE indicates that the model introduces direct bias against women that does not exist in the dataset, and that the model is more tuned to a ”classically male” symptom profile for AMI compared to clinical practice. We therefore focus on jointly reducing the NDE and NIE for AMI risk prediction.
5.4.2. Systemic Lupus Erythematosus (SLE)

### v2

@@ 4.2 Robustly estimate path-specific causal effects
Despite progress in doubly robust estimators, the dimensionality of $W$ and $Z$ remains a statistical bottleneck for finite sample estimators. We therefore assess the robustness of dimensionality reduction methods that provide the best tradeoff for robust path-specific estimation using synthetic data, where true effect sizes are known. We set the causal graph underlying the data-generating process to the SFM (Figure 2a) and initialize a different MLP to simulate each edge of the graph. We simulate synthetic data with dimensionality that closely matches the real data, with general function mappings (such as highly non-linear MLPs) to ensure that chosen dimensionality reduction will provide analogous robust estimates in real data where the ground-truth is unknown. Synthetic experiments show that doubly robust estimation alone still leads to large measurement errors in high dimensions ($\geq 500$ features) for the NDE and NIE (Figure 10). We empirically evaluate methods of dimensionality reduction to identify the method(s) that lead to the most accurate path-specific effect estimates. Doubly robust estimators are effective because they only require one set of models (either the propensity models or outcome models) to be correctly specified Li and Shen (2020). Thus, we focus on dimensionality reduction methods that seek to retain predictive power about the sensitive attribute $X$ or $Y$. For each method listed below, we test reduction to $20\%$, $40\%$, $50\%$, $60\%$, and $80\%$ of the original dimensionality $|W|$.
@@
  • Learn $Y$: We retain predictive power in $Y$ by identifying the "most important" subset of mediating features $W$ for predicting the outcome $Y$. We examine importance through LASSO logistic regression, commonly used for genetic mediation analysis Jérolon et al. (2024); Ye et al. (2021). Additionally, we train an XGBoost-based outcome model and identify features through permutation feature importance (PFI) Fisher et al. (2019), a model-agnostic measure of feature importance.
@@
4.3 Train a baseline model
The baseline model is trained on all variables in $X$, $W$, and $Z$. Our pipeline is model-agnostic, so any model architecture can be used. The evaluation of causal effects on model output requires "hard" (binary) labels; we create these labels by selecting the threshold probability that jointly optimizes sensitivity and specificity in the validation dataset.
4.4 Select target effects for which to reduce disparities
Once we select a clinical problem and establish a baseline model, we should select the causal paths to prioritize for elimination. This idea of prioritizing causal pathways is similar to the notion of "business necessity" Plečko and Bareinboim (2024); however, rather than ignoring bias through specific pathways, we recognize that regularization across all causal pathways may not be feasible while maintaining model performance ("no free lunch"). Thus, we advocate for case-specific prioritization of causal pathways that accounts for domain knowledge of health disparities and relative disparities in the data compared to baseline models. We examine 1) the relative sizes of each causal effect in the raw data and the ways in each observed causal effect may be reflective of known healthcare disparities in the specific clinical domain; 2) how causal effects may be exacerbated in the baseline model compared to the raw data; and 3) the contextual factors (either from health disparities literature or algorithmic fairness literature) that may explain the causal effects of the baseline model.
4.5 Perform and evaluate fairness interventions
@@
  • Demographic unawareness: We train models without any explicit demographic information (race and gender), including the sensitive attribute. This experiment is representative of the idea of "fairness through unawareness", a simple but widely debated debiasing method in fair ML and ML for health Höltgen and Oliver (2025); Basu (2023).
  • "Unbiased" feature selection: We identify "highly biased features" as those with a large standardized mean difference (SMD) between the specified sensitive attribute and remove these features during model training. For a specified feature count ($n$), we keep the $n$ features with the lowest SMD.
  • Greedy feature selection We greedily select features that maximize predictive performance while minimizing bias. We quantify a feature’s contribution to predictive performance using a feature importance metric and quantify its contribution to bias using SMD, as defined above. The relative contribution of SMD to the feature score is selected based on the "elbow point" that maximizes distance from the baseline importance-SMD tradeoff.
@@
  • Equalized Odds: We investigate the impact of imposing a group fairness constraint on path-specific causal fairness. Specifically, we operationalize equalized odds Hardt et al. (2016) by adding a regularization term to the loss function that penalizes differences in the true and false positive rates.
@@
We evaluate the models based on performance (AUROC) and path-specific causal effects (focusing on the "prioritized" causal pathways, as described in Section 4.4). In order to better characterize the impact of the regularized prediction model on specific subgroups of patients, we also report the Pearson Correlation between the sensitive attribute and the true outcome, as well as between the sensitive attribute and the predicted outcomes (Appendix 12-15). This allows us to compare the relationship between the sensitive attribute and the outcome in the data, baseline model, and the "fair" model. This analysis is especially important for tasks where our goal is to improve racial fairness, as we report the causal effects with respect to Black vs. White patients, and assessing the impact on other racial groups is crucial.
5 Experiments and Results
@@
We demonstrate our proposed workflow on four clinical risk prediction tasks: acute myocardial infarction (AMI), systemic lupus erythematosus (SLE), type 2 diabetes mellitus (T2DM), and schizophrenia (SCZ). We select these tasks because they span a diverse set of clinical domains and are all associated with known health disparities (Section 5.1). In each case, we select an "at-risk" cohort of individuals with related symptoms or comorbidities of the illness, and seek to predict which of these individuals will go on to develop the given disease. Our framing of tasks within the "at-risk" cohort rather than a general population increases the task difficulty because many well-established risk factors for the disease are widely present in the data and are therefore less predictive of the outcome of interest. This forces the model to rely less on risk factors that clinicians are already attuned to, thus increasing potential for clinical utility. For AMI, SLE, and T2DM, we use the phenotypes defined in a prior foundation model benchmark Pang et al. (2025) and retrain the LLAMA foundation model on CUMC-EHR. We select the LLAMA model because is a widely used architecture used both within and outside the healthcare setting Wornow et al. (2025).
5.1 Define the task and key variables
@@
We demonstrate our pipeline in two datasets representing typical structured observational health data sources: CUMC-EHR, a deidentified EHR of $\sim$5.3 million patients from a large urban academic medical center, and the deidentified Merative Multi-State Medicaid dataset (MDCD). The MDCD dataset includes all data collected from healthcare providers for medical billing for $25$ million patients from $11$ different US states. Both datasets are structured using the Observational Medical Outcomes Partnership Common Data Model Hripcsak et al. (2019), ensuring reproducibility in other compatible health datasets. In both datasets, gender is recorded as male or female and the method of capture for this variable (e.g., patient-reported, assigned at birth) is not recorded and may vary between patients. We are therefore unable to extend our analyses beyond binary gender. In CUMC-EHR, race is captured as a categorical variable: American Indian or Alaskan Native, Asian, Black, Native Hawaiian or Pacific Islander, White, Missing, and Other. In MDCD, race is captured with three categories: Black, White, and Missing. The method of capture for race is also unknown in both datasets.
5.1.1 Acute Myocardial Infarction (AMI)
@@
We predict AMI in patients with prior ischemic heart disease. We focus on mitigating gender disparities ($x_{0}=female,x_{1}=male$) due to known gender-based differences in presenting symptoms that can lead to underdiagnosis and delayed diagnosis of AMI in women (Kim et al., 2023b; Ngaruiya et al., 2024; Imboden et al., 2026). We predict AMI onset using the embeddings from a LLAMA foundation model Wornow et al. (2025) pre-trained on CUMC-EHR; the model uses RoPE embeddings to encode temporal information and is trained using next token prediction (NTP) with context length $8192$. The model embeddings represent a person’s clinical history ($W$) and can be optionally combined with $X$ and $Z$ to predict the downstream outcome. This dataset does not contain comprehensive information about visit settings; thus, we represent healthcare utilization as a single variable defined as the number of distinct days with clinical "touchpoints" (visits, diagnoses, or procedures) per year. $Z$ consists of healthcare utilization and patient race.
5.1.2 Systemic Lupus Erythematosus (SLE)
@@
We predict the first occurrence of SLE among individuals with at least one SLE-related symptom or a prescription of a drug used to treat SLE symptoms. We focus on mitigating gender disparities ($x_{0}=female,x_{1}=male$) in the case of SLE due to differing diagnosis rates ($4-22\%$ of diagnosed SLE patients are male) and significantly differing symptom profiles across sexes Tan et al. (2012); do Socorro Teixeira Moreira Almeida et al. (2011). Additionally, prior experimental work has found that clinicians faced with otherwise identical symptom profiles are less likely to diagnose men with SLE Simard et al. (2022). We use the same data source (CUMC-EHR), LLAMA model, and healthcare utilization as outlined above for AMI.
5.1.3 Type 2 Diabetes Mellitus (T2DM)
@@ 5.1.4 Schizophrenia (SCZ)
For the case of SCZ risk prediction, our "at-risk" cohort is defined as individuals with at least 3 years of observation and at least one psychosis diagnosis, and we predict schizophrenia onset one year after the first psychosis visit. We mandated three years of observation because external validation of the SCZ phenotype demonstrates that requiring three years of prior observation improves the positive predictive value from $69\%$ to $77\%$ (Finnerty et al., 2024). Black individuals are overdiagnosed with schizophrenia due to differences in the interpretation of affective symptoms Trierweiler et al. (2000); Gara et al. (2019). Additionally, rates of SCZ diagnosis following psychosis are higher for Black individuals compared to white individuals Moe et al. (2024). For these reasons, we focus on mitigating racial disparities ($x_{0}=White,x_{1}=Black$). The SCZ prediction model is trained for this single task using the MDCD data. We generated features ($W$) based on frequency of conditions, laboratory tests, procedures and duration of prescriptions. We define healthcare utilization through $10$ setting-specific frequency variables (e.g., outpatient, inpatient). $Z$ consists of healthcare utilization and patient gender.
@@
Figure 5: Error in the path-specific causal estimates for dimensionality $|W|=1,750$. Each line corresponds to a different dimensionality reduction method, with the black line corresponding to no dimensionality reduction. The NDE estimation error is reflected in part (a), and the NIE estimation error is reflected in part (b). We find that interventions that identify features important for predicting the outcome ($Y$) outperform all other methods.
@@
Figure 6: Side-by-side line plots with associated error bars. Each line is a different color, corresponding to the given legend of dimensionality reudction techniques
5.2 Robustly estimate path-specific causal effects
@@
Using synthetic data, we demonstrate that estimating causal effects in high mediator dimensions $(|W|\geq 500)$ leads to large estimation errors in the NDE and NIE (Figure 10, Table 2). We test each of the dimensionality reduction methods outlined in Section 4.2 on various mediator dimension sizes, ranging from $|W|=750$ to $|W|=1,750$. We consistently find that dimensionality reduction methods that identify the most important features for predicting $Y$, particularly PFI, are most effective at reducing both the NDE and NIE estimation error across all sample sizes (Figure 6). The spurious effect calculation is not impacted by changing the representation of $W$. These results are broadly consistent for $|W|=750$, $|W|=1,000$ and $|W|=1,500$ (Appendix 10). In the case of AMI, SLE, and T2DM, we learn the features most predictive of $Y$ through PFI (Fisher et al., 2019), as this yields the best results in the experiments where $|W|=750$ (and the clinical embeddings are all $|W|=768$). In the case of SCZ, we select highly predictive features using augmented feature occlusion (AFO), a permutation analysis method appropriate for temporal models (Tonekaboni et al., 2020).
5.3 Train a baseline model
@@
Table 1: We report the Natural Direct Effect (NDE), Natural Indirect Effect (NIE), and Spurious Effect (SE) in the dataset and the baseline model for each prediction task. We report the effect size based on the held-out test set to remain consistent with the evaluation of model regularization methods and report accompanying $95\%$ confidence intervals derived through bootstrapping.
5.4 Select target effects for which to reduce disparities
@@
5.4.1 Acute Myocardial Infarction (AMI)
@@
When examining causal effects in the data, only the NIE is significant ($0.008$ $95\%$ CI: $0.005,0.009$]); the positive value of the NIE indicates that an individual with clinical features drawn from a "male" population distribution is more likely to be diagnosed with AMI. This is consistent with findings that clinicians underdiagnose women with AMI due to a differing "typical" symptom profile (Kim et al., 2023b; Ngaruiya et al., 2024; Imboden et al., 2026). In the baseline model, both the NDE and NIE increase by an order of magnitude (NDE: $0.073$; NIE: $0.153$), while the SE remains relatively small (0.005[0.008,0.004]-0.005\ [-0.008,\ -0.004]). The increase in NDE and NIE indicates that the model introduces direct bias against women that does not exist in the dataset, and that the model is more tuned to a "classically male" symptom profile for AMI compared to clinical practice. We therefore focus on jointly reducing the NDE and NIE for AMI risk prediction.
5.4.2 Systemic Lupus Erythematosus (SLE)